# Deep Sparse Latent Feature Models for Knowledge Graph Completion

### v1

@@ 3.3 VAE decoder
We model the probability distribution $p_{\theta}$ through a decoder network $g_{\theta}$, parameterized by $\theta$. Given the latent variables $\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr}$ and $\mathbf{w}_{t}$, the decoder network generates a link $A_{hr,t}\sim p_{\theta}(A_{hr,t}|\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr}$ and $\mathbf{w}_{t})$. We first computes the Hadamard product to obtain $\mathbf{f}_{hr}$ and $\mathbf{f}_{t}$, as outlined in Equation 6. An MLP with non-linear activations is subsequently employed to transform $\mathbf{f}_{hr},\mathbf{f}_{t}$ into $\mathbf{g}_{hr},\mathbf{g}_{t}$, respectively.
@@ 3.4 Inference
$\mathcal{L}$ $=\sum_{(h,r)\in\mathcal{Q}}\left\{D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{v}_{hr})||p_{\theta}(\mathbf{v}_{hr})\right]+D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{hr})||p_{\theta}(\mathbf{z}_{hr}|\mathbf{v}_{hr})\right]+D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{w}_{hr})||p_{\theta}(\mathbf{w}_{hr})\right]\right\}$
$~{}~{}+\sum_{t\in\mathcal{E}}\left\{D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{v}_{t})||p_{\theta}(\mathbf{v}_{t})\right]+D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{z}_{t})||p_{\theta}(\mathbf{z}_{t}|\mathbf{v}_{t})\right]+D_{\mathrm{KL}}\left[q_{\phi}(\mathbf{w}_{t})||p_{\theta}(\mathbf{w}_{t})\right]\right\}$
$~{}~{}-\sum_{(h,r)\in\mathcal{Q}}\mathbb{E}_{q}\left[\log p_{\theta}(\mathbf{x}_{hr}|\mathbf{z}_{hr},\mathbf{w}_{hr})\right]-\sum_{t\in\mathcal{E}}\mathbb{E}_{q}\left[\log p_{\theta}(\mathbf{x}_{t}|\mathbf{z}_{t},\mathbf{w}_{t})\right]$
$~{}~{}-\sum_{(h,r)\in\mathcal{Q}}\sum_{t\in\mathcal{E}}\mathbb{E}_{q}\left[\log p_{\theta}(A_{hr,t}|\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr},\mathbf{w}_{t})\right]$ (17)
@@
To further enhance KGC performance, we express the triple completion term $\log p_{\theta}(A_{hr,t}|\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr},\mathbf{w}_{t})$ as a contrastive loss. The contrastive framework is renowned for its capacity to learn expressive representations, as it aims to maximize the mutual information between the inputs and the outputs (Ben-Shaul et al., 2023; Hjelm et al., 2018; Gutmann & Hyvärinen, 2012). Specifically, we utilize the supervised contrastive loss (Li et al., 2023a; Khosla et al., 2020):
$\log p_{\theta}(A_{hr,t}|\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr},\mathbf{w}_{t})=\frac{1}{|\mathcal{N}^{+}|}\sum_{t\in\mathcal{N}^{+}}\log\frac{e^{(S(\mathbf{g}_{hr},\mathbf{g}_{t})-\gamma)/\tau}}{e^{(S(\mathbf{g}_{hr},\mathbf{g}_{t})-\gamma)/\tau}+\sum_{t^{\prime}\in\mathcal{N}^{-}}{e^{(S(\mathbf{g}_{hr},\mathbf{g}_{t^{\prime}})-\gamma)/\tau}}}$ (18)
where $\mathcal{N}^{+}$ represents the set of positive entities of the query $(h,r,?)$, and $\mathcal{N}^{-}$ denotes the set of negative samples, encompassing all other entities within the same batch (Chen et al., 2020). The variable $\gamma$ denotes the additive margin, $\tau$ the temperature and $S(\mathbf{g}_{hr},\mathbf{g}_{t})=\cos(\mathbf{g}_{hr},\mathbf{g}_{t})=\mathbf{g}_{hr}^{\top}\mathbf{g}_{t}/(||\mathbf{g}_{hr}||\cdot||\mathbf{g}_{t}||)\in[-1,1]$ the cosine similarity score function.
We then optimize the objective using Stochastic Gradient Variational Bayes (SGVB) and mini-batch gradient descent (Kingma & Welling, 2013). Given a batch of triples $B\subset\mathcal{G}$, and let the decoded representations $\mathbf{g}_{hr},\mathbf{g}_{t}\in\mathbb{R}^{D}$, the computation of $\mathcal{L}$ requires time $\mathcal{O}(|B|\cdot(C_{\text{KL}}+C_{\text{Recon}}+C_{\text{Comp}}))$ and space $\mathcal{O}(|B|\cdot D+|B|\cdot K)$, where $C_{\text{KL}}$, $C_{\text{Recon}}$ and $C_{\text{Comp}}$ denotes the complexity of evaluating the KL divergence, reconstruction and triple completion terms in the ELBO, respectively. Please refer to Appendix A for additional proofs and computation details.
@@ 4.1 Experiment settings
Datasets. To evaluate our method for filling in missing triples in KGs, we selected benchmark datasets ranging from moderate-sized (about 93k triples) to large-scale (around 20 million triples) for the KGC task. These include WN18RR, FB15k-237 (Toutanova et al., 2015), and Wikidata5M (Wang et al., 2021b). Originally introduced by Bordes et al. (2013), the WN18 and FB15k datasets were later refined to WN18RR and FB15k-237 following studies (Toutanova et al., 2015; Dettmers et al., 2018) that revealed test leakage issues. Textual data comes from KG-BERT (Yao et al., 2019), while Wikidata5M (Wang et al., 2021b) is a large-scale KG merging Wikidata and Wikipedia, with textual descriptions for each entity.
Evaluation metrics. In our approach, for each query $(h,r,?)$, a score is calculated for each entity and the rank of the correct answer is determined. We report the Mean Reciprocal Rank (MRR) and Hit@$k$ metrics under the filtered protocol (Bordes et al., 2013). For each triple $(h,r,t)$, we construct a forward query $(h,r,?)$ with $t$ as the answer, along with a backward query (?$,r^{-1},t)$ for data augmentation. Here, $r^{-1}$ denotes the inverse of the relation $r$, as sourced from Li et al. (2023a). The averaged results of the forward and backward metrics are reported in our experimental evaluations.
Baselines. We conduct comprehensive experiments to evaluate the performance of our model against a variety of KGC models, encompassing rule-based, embedding-based and text-based KGC approaches.
Implementation details. To ensure a fair comparison with existing approaches, we maintain the same primary hyperparameters. Specifically, the BERT encoders are initialized with pre-trained weights from "bert-base-uncased". We use a batch size of 1024 with 4 Quadro RTX 8000 GPUs, although a larger batch size is reasonably expected to provide better performance under the contrastive framework. The maximum number of communities $K$ is consistently set to 128 for all datasets. In the case of the WN18RR and FB15k-237 datasets, we utilize in-batch negative sampling, whereas for the Wikidata5M dataset, we adopt an additional self-negative sampling strategy to ensure fair comparison with SimKGC (Wang et al., 2022).
Detailed information regarding the experimental setup can be found in Appendix B.1.
@@ 4.2 Main results
Table 1: Knowledge graph completion results for the WN18RR and FB15k-237 datasets.
Method WN18RR FB15k-237
MRR Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10
Rule-based Methods
NeuralLP 38.1 36.8 38.6 40.8 23.7 17.3 25.9 36.1
DRUM 38.2 36.9 38.8 41.0 23.8 17.4 26.1 36.4
LERP 62.2 59.3 63.4 68.2 - - - -
Embedding-based Methods
TransE 24.3 4.3 44.1 53.2 27.9 19.8 37.6 44.1
DistMult 44.4 41.2 47.0 50.4 28.1 19.9 30.1 44.6
R-GCN 12.3 8.0 13.7 20.7 16.4 10.0 18.1 30.0
RotatE 47.6 42.8 49.2 57.1 33.8 24.1 37.5 53.3
TuckER 47.0 44.3 48.2 52.6 35.8 26.6 39.4 54.4
HittER 50.3 46.2 51.6 58.4 37.3 27.9 40.9 55.8
N-Former 48.6 44.3 50.1 57.8 37.2 27.7 41.2 55.6
KRACL 52.7 48.2 54.7 61.3 36.0 26.6 39.5 54.8
Text-based Methods
KG-BERT 21.6 4.1 30.2 52.4 - - - 42.0
MTL-KGC 33.1 20.3 38.3 59.7 26.7 17.2 29.8 45.8
StAR 40.1 24.3 49.1 70.9 29.6 20.5 32.2 48.2
SimKGC 66.6 58.7 71.7 80.0 33.6 24.9 36.2 51.1
KG-S2S 57.4 53.1 59.5 66.1 33.6 25.7 37.3 49.8
GHN 67.8 59.6 71.9 82.1 33.9 25.1 36.4 51.8
DSLFM-KGC 70.4 63.1 74.8 84.2 35.5 26.4 38.9 53.7
Due to the stochastic nature of our model, we perform five independent experiments with different random seeds and report the average metrics. Table 2 presents the results for the Wikidata5M dataset, while Table 1 for the WN18RR and FB15k-237 datasets. Hit@$k$ is expressed as a percentage. The best performance for each metric in each dataset is highlighted in bold, and the top metrics across categories are underlined.
Table 2: KGC results for the Wikidata5M datasets.
@@
DKRL 23.1 5.9 32.0 54.6
KEPLER 40.2 22.2 51.4 73.0
BLP-ComplEx 48.9 26,2 66.4 87.7
BLP-SimplE 49.3 28.9 63.9 86.6
SimKGC 71.3 60.7 78.7 91.3
@@
The most substantial improvement is seen on the Wikidata5M dataset, where our model shows a 5.0% increase in MRR (from 71.3% to 76.3%) and a 6.5% increase in Hit@1 (from 60.7% to 67.2%) compared to SimKGC. Similar improvements are observed on the WN18RR dataset, where DSLFM-KGC surpasses the second-best model (GHN) across all metrics, with enhancements ranging from 1.9% to 3.5% in MRR and Hit@k, demonstrating its strong predictive capability. On the FB15k-237 dataset, while our model falls short of embedding-based models, it still outperforms rule-based and text-based methods, narrowing the gap between text-based and embedding-based approaches by approximately 2-3 percentage points.
To clarify the results obtained from the WN18RR and FB15k-237 datasets, we perform a detailed analysis of the underlying KGs. First, we assess the topological structure of each KG by calculating the average degree $M/N$, where $M$ and $N$ represent the number of edges and nodes, respectively. The FB15k-237 dataset exhibits a denser structure, with an average degree of 21.3, compared to 2.27 for WN18RR. Second, we examine the topological structures of both datasets. In FB15k-237, relationships show a high degree of correlation (e.g., ‘award nominee’, ‘nominee of award’), resulting in a densely interconnected structure with a less pronounced clustering pattern. Finally, we carry out in-depth ablation studies to further examine the challenges our model experiences when capturing latent community structures from the FB15k-237 dataset, as discussed in the following section.
@@ 4.3 Ablation results
We conduct diverse ablation experiments to investigate into how key components of our model impact KGC performance.
Stick-breaking prior. We conduct KGC experiments with $\alpha_{\text{qry}}$ and $\alpha_{\text{ans}}$ chosen from the grid $\{80,90,100\}\times\{10,20,\ldots,100\}$, while keeping all other hyperparameters fixed. Table 8 reports the mean and standard deviation of these 30 results for each dataset. The minimal variation in performance with different $\alpha_{\text{qry}}$ and $\alpha_{\text{ans}}$ values, as seen in Table 8, highlights the robustness of our model under diverse prior settings.
@@
As discussed in Section 4.2, the denser topology of the FB15k-237 dataset makes it more difficult to capture community structures. To gain further insights, we compute the average number of activated communities (the number of non-zero entries in $Z_{\text{ans}}$ divided by the number of entities $|\mathcal{E}|$) and present the trend across different $\alpha_{\text{ans}}$ values in Figure 3. Clearly, for identical $\alpha_{\text{ans}}$ values, FB15k-237 exhibits significantly fewer latent communities than WN18RR, with the disparity increasing as $\alpha_{\text{ans}}$ rises. This indicates the greater density and less pronounced clustering structure of the FB15k-237 dataset.
Table 3: Performance of DSLFM-KGC on the WN18RR and FB15k-237 datasets w/ different latent structures.
Method WN18RR FB15k-237
Hit@1 Hit@10 Epochs Hit@1 Hit@10 Epochs
Ours 63.1 84.2 65 26.4 53.7 15
VAE 60.9 82.4 55 25.6 52.0 10
AE 59.0 82.0 50 25.1 52.3 10
Dose the sparse latent structure makes a difference? To assess this, we replace our encoder with one that generates an approximate standard Gaussian distribution, as used in the vanilla VAE (Kingma & Welling, 2013). Additionally, we evaluate a pure autoencoder (AE), which assumes no probabilistic distribution for the latent variables. The testing performance and the training convergence epochs (based on the best validation metric) for the WN18RR and FB15k-237 datasets are shown in Table 3.
The results in Table 3 show that integrating latent structure substantially enhance KGC performance on the WN18RR dataset. However, the FB15k-237 dataset witnesses only modest improvements, illustrating the challenges in modeling its latent community structure. Furthermore, the increased complexity of the latent structure negatively impacted the convergence rate, as evidenced by the longer training epochs. Future studies to enhance KGC accuracy on dense-connected KGs and improve training efficiency are necessary.
(a) WN18RR
(b) FB15k-237

### v2

@@ 3.3 VAE decoder
We model the probability distribution $p_{\theta}$ through a decoder network $g_{\theta}$, parameterized by $\theta$. Given the latent variables $\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr}$ and $\mathbf{w}_{t}$, the decoder network generates a link $A_{hr,t}\sim p_{\theta}(A_{hr,t}|\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr},\mathbf{w}_{t})$. We first computes the Hadamard product to obtain $\mathbf{f}_{hr}$ and $\mathbf{f}_{t}$, as outlined in Equation 6. An MLP with non-linear activations is subsequently employed to transform $\mathbf{f}_{hr},\mathbf{f}_{t}$ into $\mathbf{g}_{hr},\mathbf{g}_{t}$, respectively.
@@ 3.4 Inference
$\mathcal{L}$ $=D_{\mathrm{KL}}[q_{\phi}(\mathcal{H})||p_{\theta}(\mathcal{H})]-\mathbb{E}_{q}\left[\log p_{\theta}(\mathcal{O}|\mathcal{H})\right]$
$=D_{\mathrm{KL}}[q_{\phi}(\mathcal{H}_{\text{qry}})||p_{\theta}(\mathcal{H}_{\text{qry}})]+D_{\mathrm{KL}}[q_{\phi}(\mathcal{H}_{\text{ans}})||p_{\theta}(\mathcal{H}_{\text{ans}})]$
$~{}-\mathbb{E}_{q}\left[\log p_{\theta}(\mathcal{Q}|\mathcal{H}_{\text{qry}})\right]-\mathbb{E}_{q}\left[\log p_{\theta}(\mathcal{E}|\mathcal{H}_{\text{ans}})\right]-\mathbb{E}_{q}\left[\log p_{\theta}(A|\mathcal{H})\right]$ (18)
@@
To further enhance KGC performance, we express the triple completion term $\log p_{\theta}(A|\mathcal{H})$ as a contrastive loss for its renowned capacity to learn expressive representations, which aims to maximize the mutual information between the inputs and the outputs (ben2023reverse, ; hjelm2018learning, ; gutmann2012noise, ). Specifically, given the set of latent variables for a specific query and answer, we utilize the supervised contrastive loss (li2023kermit, ; khosla2020supervised, ):
$\log p_{\theta}(A_{hr,t}|\mathbf{z}_{hr},\mathbf{z}_{t},\mathbf{w}_{hr},\mathbf{w}_{t})=\frac{1}{|\mathcal{N}^{+}|}\sum_{t\in\mathcal{N}^{+}}\log\frac{e^{S(\mathbf{g}_{hr},\mathbf{g}_{t})}}{e^{S(\mathbf{g}_{hr},\mathbf{g}_{t})}+\sum_{t^{\prime}\in\mathcal{N}^{-}}{e^{S(\mathbf{g}_{hr},\mathbf{g}_{t})}}}$ (19)
where $\mathcal{N}^{+}$ represents the set of positive entities of the query $(h,r,?)$, and $\mathcal{N}^{-}$ denotes the set of negative samples, encompassing all other entities within the same batch (pmlr-v119-chen20j, ). $S(\mathbf{g}_{hr},\mathbf{g}_{t})=(\cos(\mathbf{g}_{hr},\mathbf{g}_{t})-\gamma)/\tau$ is the cosine similarity score function with additive margin $\gamma$ and temperature $\tau$, following the work of (kumar2018mises, ).
We then optimize the objective using stochastic gradient variational Bayes (SGVB) and mini-batch gradient descent (kingma2013auto, ). Given a batch of triples $B\subset\mathcal{G}$, and let the decoded representations $\mathbf{g}_{hr},\mathbf{g}_{t}\in\mathbb{R}^{D}$, the computation of $\mathcal{L}$ requires time $\mathcal{O}(|B|\cdot(C_{\text{KL}}+C_{\text{Recon}}+C_{\text{Comp}}))$ and space $\mathcal{O}(|B|\cdot D+|B|\cdot K)$, where $C_{\text{KL}}$, $C_{\text{Recon}}$ and $C_{\text{Comp}}$ denotes the complexity of evaluating the KL divergence, reconstruction and triple completion terms in the ELBO, respectively. Please refer to Appendix B for additional proofs and computation details.
@@ 4 Experiment
Table 1: KGC results for the WN18RR, FB15k-237 and UMLS datasets.
Method WN18RR FB15k-237 UMLS
MRR Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10 MRR Hit@1 Hit@3 Hit@10
Embedding-based Methods
TransE bordes2013translating 24.3 4.3 44.1 53.2 27.9 19.8 37.6 44.1 61.5 39.1 - 98.9
DistMult yang2014embedding 44.4 41.2 47.0 50.4 28.1 19.9 30.1 44.6 39.1 25.6 44.5 66.9
RotatE sun2018rotate 47.6 42.8 49.2 57.1 33.8 24.1 37.5 53.3 74.4 63.6 82.2 93.9
ConvE dettmers2018convolutional 43.0 40.0 44.0 52.0 32.5 23.7 35.6 50.1 83.6 76.4 96.0 99.0
TuckER balavzevic2019tucker 47.0 44.3 48.2 52.6 35.8 26.6 39.4 54.4 73.2 62.5 81.2 90.0
HittER chen-etal-2021-hitter 50.3 46.2 51.6 58.4 37.3 27.9 40.9 55.8 - - - -
KRACL tan2023kracl 52.7 48.2 54.7 61.3 36.0 26.6 39.5 54.8 90.4 83.1 - 99.5
GNN-based Methods
RGCN schlichtkrull2018modeling 12.3 8.0 13.7 20.7 16.4 10.0 18.1 30.0 48.1 31.8 - 83.5
CompGCN vashishth2019composition 47.2 43.7 48.5 54.0 35.5 26.4 39.0 53.6 - - - -
KGGAT nathani2019learning 46.4 42.6 47.9 53.9 35.0 26.0 38.5 53.1 - - - -
Text-based Methods
KG-BERT yao2019kg 21.6 4.1 30.2 52.4 - - - 42.0 - - - 99.0
MTL-KGC (kim-etal-2020-multi, ) 33.1 20.3 38.3 59.7 26.7 17.2 29.8 45.8 - - - -
StAR (wang2021structure, ) 40.1 24.3 49.1 70.9 29.6 20.5 32.2 48.2 - - - 99.1
SimKGC wang2022simkgc 66.6 58.7 71.7 80.0 33.6 24.9 36.2 51.1 79.4 70.4 86.5 94.4
KG-S2S (chen2022knowledge, ) 57.4 53.1 59.5 66.1 33.6 25.7 37.3 49.8 - - - -
GHN (qiao-etal-2023-improving, ) 67.8 59.6 71.9 82.1 33.9 25.1 36.4 51.8 - - - -
LLM-based Methods
CP-KGC yang2024enhancing 67.3 59.9 72.1 80.4 33.8 25.1 36.5 51.6 - - - -
KICGPT wei2024kicgpt 54.9 47.4 58.5 64.1 41.2 32.7 44.8 55.4 - - - -
DSLFM-KGC (ours) 70.4 63.1 74.8 84.2 35.5 26.4 38.9 53.7 92.4 86.4 98.7 99.6
@@ 4.1 Experiment settings
Datasets. To evaluate our method for filling in missing triples in KGs, we selected benchmark datasets ranging from small-sized (about 6k triples) to large-scale (around 20 million triples) for the KGC task. These include UMLS kok2007statistical , WN18RR, FB15k-237 (toutanova2015representing, ), and Wikidata5M-Ind (wang2021kepler, ).
Evaluation metrics. In our approach, for each query $(h,r,?)$, a score is calculated for each entity and the rank of the correct answer is determined. We report the Mean Reciprocal Rank (MRR) and Hit@$k$ metrics under the filtered protocol (bordes2013translating, ). For each triple $(h,r,t)$, we construct a forward query $(h,r,?)$ with $t$ as the answer, along with a backward query (?$,r^{-1},t)$ for data augmentation. Here, $r^{-1}$ denotes the inverse of the relation $r$, as sourced from li2023kermit . The averaged results of the forward and backward metrics are reported in our experimental evaluations.
Baselines. We conduct comprehensive experiments to evaluate the performance of our model against a variety of KGC models, encompassing embedding-based, GNN-based, text-based and LLM-based KGC approaches.
Implementation details. To ensure a fair comparison with existing approaches, we maintain the same primary hyperparameters. Specifically, the BERT encoders are initialized with pre-trained weights from "bert-base-uncased". We use a batch size of 1024 with 4 Quadro RTX 8000 GPUs, although a larger batch size is reasonably expected to provide better performance under the contrastive framework. The maximum number of communities $K$ is consistently set to 128 for all datasets. In the case of the WN18RR, FB15k-237 and UMLS datasets, we utilize in-batch negative sampling, whereas for the Wikidata5M-Ind dataset, we adopt an additional self-negative sampling strategy to ensure fair comparison with SimKGC (wang2022simkgc, ).
Detailed information regarding the experimental setup can be found in Appendix C.
@@ 4.2 Main results
Figure 2: KGC results for the Wikidata5M-Ind dataset.
@@
DKPL (xie2016representation, ) 23.1 5.9 32.0 54.6
KEPLER (wang2021kepler, ) 40.2 22.2 51.4 73.0
BLP-ComplEx daza2021inductive 48.9 26,2 66.4 87.7
BLP-SimplE daza2021inductive 49.3 28.9 63.9 86.6
SimKGC wang2022simkgc 71.3 60.7 78.7 91.3
@@
Owing to the stochastic nature of our model, we conduct five independent experiments with distinct random seeds and report the average metrics. Table 1 presents the results for the WN18RR and FB15k-237 datasets, while Table 2 presents the results for the Wikidata5M-Ind dataset. Hit@k is expressed as a percentage. The highest performance for each metric on each dataset is indicated in bold, and the top-performing metrics across categories are underlined.
The most substantial improvement is observed on the Wikidata5M-Ind dataset, where our model exhibits a 5.0% increase in MRR (from 71.3% to 76.3%) and a 6.5% increase in Hit@1 (from 60.7% to 67.2%) compared to SimKGC. Similar improvements are observed on the WN18RR dataset, where DSLFM-KGC surpasses the second-best model (GHN) across all metrics, with enhancements ranging from 1.9% to 3.5% in MRR and Hit@k, demonstrating its strong predictive capability. On the FB15k-237 dataset, while our model falls short of embedding-based models and KICGPT wei2024kicgpt , it still outperforms GNN-based and text-based methods, narrowing the gap between text-based and embedding-based approaches by approximately 2-3 percentage points. The results demonstrate both the effectiveness and the scalability of our model.
To elucidate the results obtained from the WN18RR and FB15k-237 datasets, we perform a detailed analysis of the underlying KGs. i) First, we assess the topological structure of each KG by calculating the average degree $M/N$, where $M$ and $N$ represent the number of edges and nodes, respectively. The FB15k-237 dataset exhibits a denser structure, with an average degree of 21.3, compared to 2.27 for WN18RR. ii) Second, we measure the modularity of the KGs and present the results in Appendix D.2. The significantly lower modularity score for FB15k-237 (0.074) quantitatively confirms our hypothesis that FB15k-237 exhibits much less pronounced clustering patterns. iii) Finally, we conduct in-depth ablation and analysis studies to further investigate the challenges our model encounters when capturing latent community structures from the FB15k-237 dataset, as discussed in subsequent sections.
@@ 4.3 Ablation results
We conduct a series of ablation experiments to investigate the impact of key model components on KGC performance.
@@
Robustness w.r.t. hyperparameter settings. We conduct KGC experiments with $\alpha_{\text{qry}}$ and $\alpha_{\text{ans}}$ selected from the grid $\{80,90,100\}\times\{10,20,\ldots,100\}$, while maintaining all other hyperparameters fixed. Table 6 reports the mean and standard deviation of these 30 results for each dataset. The minimal variability observed across different $\alpha_{\text{qry}}$ and $\alpha_{\text{ans}}$ values, as seen in Table 6, underscores the robustness of our model to varying prior settings, confirming that our model maintains a favorable balance between completing the KGC task and determining the number of clusters. Further analysis of the model’s robustness to other hyperparameters is provided in Appendix D.1.
As discussed in Section 4.2, the denser topology of the FB15k-237 dataset presents a greater challenge for capturing community structures. To gain further insight, we compute the average number of activated communities (the number of non-zero entries in $Z_{\text{ans}}$ divided by the number of entities $|\mathcal{E}|$) and present the trend across varying $\alpha_{\text{ans}}$ values in Figure 3. Clearly, for identical $\alpha_{\text{ans}}$ values, FB15k-237 exhibits significantly fewer latent communities than WN18RR, with the disparity increasing as $\alpha_{\text{ans}}$ rises. This indicates the greater density and less pronounced clustering structure of the FB15k-237 dataset.
Table 2: Performance of DSLFM-KGC on the WN18RR, FB15k-237 and UMLS datasets w/ different latent structures.
Method WN18RR FB15k-237 UMLS
Hit@1 Hit@10 Epochs Hit@1 Hit@10 Epochs Hit@1 Hit@10 Epochs
Ours 63.1 84.2 65 26.4 53.7 15 86.4 99.6 25
VAE 59.1 81.0 55 25.4 51.5 10 71.5 89.0 23
AE 57.9 80.0 50 25.2 52.0 10 69.6 92.7 19
SimKGC 58.7 80.8 50 24.9 51.1 10 70.4 94.4 20
Figure 4: Comparison of DSLFM-KGC and SimKGC performance on the WN18RR dataset w.r.t. neighborhoods at varying geodesic distances, where ‘5+’ denotes a distance greater than 5. If no path exists between two entities, the distance is considered infinite.
Does modeling sparse community structure really benefit KGC?. To evaluate the impact of the sparse latent structure, we replace only our posterior encoder with one that generates an approximate standard Gaussian distribution, as employed in the vanilla VAE (kingma2013auto, ), while keeping the BERT and MLP components unchanged. Additionally, we evaluate a pure autoencoder (AE), which does not assume a probabilistic distribution for the latent variables. The testing performance and the training convergence epochs (based on the best validation metric) for the WN18RR, FB15k-237 and UMLS datasets are presented in Table 2.
Table 2 demonstrates that incorporating latent structure significantly improves KGC performance on the WN18RR and UMLS datasets. Furthermore, the significant performance difference observed between our model and its variants, coupled with the close results obtained by the VAE, AE, and SimKGC models, indicates that the enhancement in KGC performance arises from the sparse community modeling itself, rather than solely from the increased model complexity. However, on the FB15k-237 dataset, improvements are more modest, suggesting challenges in accurately capturing its underlying latent community structure. Moreover, the increased complexity introduced by the latent structure model negatively impacted convergence speed, as indicated by the longer training epochs. Future research should focus on improving KGC accuracy on densely connected KGs while simultaneously enhancing training efficiency.